)Assume that only two countries, A and B, exist Consider the following data Factor Endowments 45 15 Labor 10 Cepited Stck If good S is capital intensive, then following the Heckscher-Ohlin Theory A) country A will export good S. B) country B will export good S C) both countries will export good S D) trade will not occur between these two countries. E) Insufficient information is given. 8) Continuing from Question #7, you now are told that the labor unions representing the workers in each of the two respective countries are considering lobbying agai Note that workers income is derived solely from wages. inst the opening of international trade between these two countries. A) This would be a good decision from the viewpoint of workers in Country B) This would be a misguided decision from the viewpoint of the workers in C) This would be a good decision from the viewpoint of workers in Country D) This would be a good decision for both unions since trade hurts workers E) None of the above. A since intenational trade helps the capitalists and hurts the workers. both countries, since trade is always better than autarky B since international trade helps the owners of capital and hurts the workers wherever they live. 9) Export-biased growth in Country H will A) improve the terms of trade of Country F B) decrease economic welfare in Country H C) improve the terms of trade of Country H D) worsen the terms of trade of Country F (the trade partner). E) trigger anti-bias regulations of the WTO 10) If the U.S. (a large country) imposes a tariff on its imported good, this will tend to A) improve the terms of trade of the United States. B) cause a deterioration of U.S. terms of trade. C) raise the world price of the good imported by the United States. D) improve the terms of trade of all countries. E) have no effect on terms of trade. 11) Internal economies of scale A) may be associated with a perfectly competitive industry. B) cannot be associated with a perfectly competitive industry C) are associated only with sophisticated products such as aircraft. D) cannot form the basis for international trade E) None of the above. 12) The larger the number of firms in a monopolistic competition situation, A) the larger are that countrys exports. B) the lower is the price charged. C) the higher is the price charged. D) the fewer varieties are sold. E) None of the above.

7 to 12 answers I need please this is MCQ
Answer #1

Country A

Labor force = 45

Capital stock = 15

Calculate the capital to labor ratio of Country A -

Capital to Labor ratio = Capital stock/Labor force = 15/45 = 0.33

The capital to labor ratio of country A is 0.33

Country B

Labor force = 20

Capital stock = 10

Calculate the capital to labor ratio of Country B -

Capital to Labor ratio = Capital stock/Labor force = 10/20 = 0.50

The capital to labor ratio of country B is 0.50

It can be seen that the capital to labor ratio of Country B is greater than that of the Country A.

So,

Country B is capital intensive country.

Hecksher-Ohlin theory states that a country exports that good which it can produce using its abundant factor.

Since, country B is capital intensive relative to country A, it will export the capital intensive good.

Thus, if good S is capital intensive then Country B will export good S.

Hence, the correct answer is the option (B).